this one is called order of operation, graphing, and problem solving
I got one problem that has me stumped can you help me solve.
1.Evaluate: 3 time/16 divide by 68-3 times 5 time or it reads 3 time /16 +68-3 time 5 time 2. so can you help me with this problem


Answer by rfer(16322)

I will try, you have the two problems written differently so I am guessing some.
3*16/68-3*5=
48/68-15=
0.7-15=-14.3
That doesn't make much since to me.
-----------------------
3*16+68-3*5*2=
48+16-30=34
Second one makes better since.
Take another look at the problem and if you want send it to me on my e-mail.
Bob